Meeting-07282022

From Cuawiki
Jump to: navigation, search

General Information

  • Logbook
https://logbooks.jlab.org/book/nps
  • If you are not already subscribed please sign up to nps and nps_daq at:
https://mailman.jlab.org/mailman/listinfo/nps
https://mailman.jlab.org/mailman/listinfo/nps_daq
  • Slack
https://nps-jo93164.slack.com
  • NPS on CUE system
Brad's Hall A/C computing tricks and tips
/work/hallc/nps for development code, etc (NB not backed up)
/volatile/hallc/nps for large data file (write to tape if anything you want to keep: /mss/hallc/c-nps/analysis)
/mss/hallc/c-nps/raw is where raw data will go.
  • hcana and replay set up instructions (for pionLT run)
https://wiki.jlab.org/cuawiki/index.php/Meeting-04072022

Zoom link

https://jlab-org.zoomgov.com/j/16162588644?pwd=eklwZko1TDZkam1Ba3dSYVkyVlk4dz09

Agenda

  • DAQ and firmware
  • Slow controls
  • Software
  • Other business

Software task list (draft)

  • Near-term priorities (for summer tests)
    • hcana FADC mode 10 code (Mark, David) -- DONE
    • collect everything together on GitHub (David)
    • hcana NPS class, replay script and database file (David,??)
    • online display and cosmic calibration scripts (Julie, Pramita, Alexandre, Zheng)
    • code to handle VTP data in hcana (David)
  • Medium-term (summer - fall)
    • data unblocking (David, Brad, Bryan)
    • geant4 simulation (Marie, Deb)
    • online interface with ET system (Paul)
  • Longer term
    • offline reconstruction and clustering (Steve, Carlos Y, Julie, Pramita)
    • offline calibration scripts (??)

DAQ update (from Brad)

  • Bryan Moffit and Brad are setting up one crate with FADCs. First step will be taking pulser driven FADC pedestal data and setting up decoding in hcana (David H). Ben and Bryan are updating the firmware to be compatible with the VTP trigger.
  • High frequency (250 MHz, 200 mV) noise clear on the scope. Fernando and team are looking into it and have been able to re-create it on the bench. They suspect it is being generated in the PMT bases and that additional grounding / EM shielding will be needed.
  • There is also an overcurrent issue with one of the LV boards, which Fernando is looking at.
  • Cable bundling is ongoing in the ESB.
  • Brad asked Alex about the FADC modules that are on loan and when they could be returned.
  • There was a question to Carlos regarding how to set up the built-in trigger scintillator paddles. There are external paddles available for now while a solution is found (need a bypass LV board).

Software update

Next meeting

Thursday 18th August at 8.30am EST